Evaluate the d^3f/dx^3 of f(x)=sin(x)cos(x)
1
Expert's answer
2019-05-14T12:34:30-0400

"f(x)=sin(x)cos(x)=\frac{1}{2}sin(2x)\\
\frac{df}{dx}=cos(2x)\\
\frac{d^2f}{dx^2}=-2sin(2x)\\
\frac{d^3f}{dx^3}=-4cos(2x)"